OpenStudy (anonymous):

Whats x in 76+90=32(75)-x

OpenStudy (sleepyjess):

Do you know what your first step would be?

OpenStudy (anonymous):

76+90=32(75)-x Since x is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation. 32(75)-x=76+90 Add 90 to 76 to get 166. 32(75)-x=166 Multiply 32 by each term inside the parentheses. 2400-x=166 Since 2400 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 2400 from both sides. -x=-2400+166 Add 166 to -2400 to get -2234. -x=-2234 Multiply each term in the equation by -1. -x*-1=-2234*-1 Multiply -x by -1 to get x. x=-2234*-1 Multiply -2234 by -1 to get 2234. x=2234

OpenStudy (anonymous):

32 x 75

OpenStudy (sleepyjess):

Nevermind... lol

OpenStudy (anonymous):

Lol thanks jess anyways

OpenStudy (anonymous):

xD sorry

OpenStudy (sleepyjess):

@thewonderfuladele ... lol

OpenStudy (tkhunny):

"76+90=32(75)-x" There is not now, nor will there ever be a good answer for "the" first step. You should do whatever comes to your mind that might lead to a solution. Think about it. Ponder it. What makes the expressions complicated? What can you do to simplify the expressions and make more sense of the whole equation?

OpenStudy (sleepyjess):

@tkhunny There is a first step... your first step may not be the same as my first step, but there is always a first step. I was asking elderkid if they knew what the first step would be... was trying to find out where they didn't understand it and walk through it with them. :-)

OpenStudy (tkhunny):

Substantially disagree, but the discussion is of no consequence on this thread. Please continue your usual excellent and patient work. :-)

OpenStudy (sleepyjess):

@tkhunny ... lol... ok

OpenStudy (anonymous):

wow nyz questtıon